OpenStudy (anonymous):

solve the equation -6+8|4x-2|=58

OpenStudy (anonymous):

-6+8|4x-2|=58 8|4x-2| = 58+6 |4x-2| = 64/8 |4x-2| = 8 1. 4x-2 = 8 AND (i think) 2. 4x-2 = -8 1. 4x-2=8 4x = 10 x = 5/2 2. 4x-2 = -8 4x = -6 x = -3/2 x = 5/2 and -3/2